a question on the homework i have says: Let the sum of an arithmetic sequence of eight terms be represented by the letter S. If the first term is "a" and the eighth term is "a/2" express "a" as a function of S.
Answer by stanbon(75887)

a question on the homework i have says: Let the sum of an arithmetic sequence of eight terms be represented by the letter S.
-------------------------
If the first term is "a" and the eighth term is "a/2" express "a" as a function of S.
-------------
a(8) = a + 7d = a/2
7d = -a/2
----------
S(8) = (8/2)(a+a+7d) = 4(2a+7d)
---
S(8) = 4(2a-(a/2))
---
S(8) = 4(3/2)a)
S(8) = 6a
a = (1/6)*S(8)
=====================
Cheers,
Stan H.